lucille recieves an annual salary of $37,500 based on a 37.5 hour workweek. What are her gross earnings for a two-week pay period in which she works 9 hours of overtime at 1.5 times her regular rate of pay?
Answer by checkley71(8403)

37500/(52*37.5)=37500/1950=$19.23 PER HOUR IS HER RATE. **([52*37.5=1,950])**
19.23(37.5*2)+19.23*9*1.5=19.23*75+259.61=1,442.25+259.61=$1,701.86 IS HER PAY
FOR TWO WEEKS WORK WITH 9 HOURS OF OVERTIME @ 1.5 TIMES HER BASE SALARY.